# A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Code Smell Detection

### v2

@@ Abstract.
Code smells, which are suboptimal coding practices that can potentially lead to defects or maintenance issues, can negatively impact the quality of software systems. Most existing code smell detection methods rely on heuristics-based or machine learning (ML) and deep learning (DL)-based techniques. However, these techniques have several drawbacks (e.g., unsatisfactory performance). Large Language Models (LLMs) have garnered significant attention in the Software Engineering (SE) field, achieving state-of-the-art performance across a wide range of SE tasks. Parameter-Efficient Fine-Tuning (PEFT) methods, which are commonly used to adapt LLMs to specific tasks with fewer parameters and reduced computational resources, have emerged as a promising approach for enhancing the performance of LLMs in various SE tasks. However, LLMs have not yet been explored for code smell detection, and their effectiveness for this task remains unclear. Furthermore, no comprehensive investigation has been conducted on the efficiency of PEFT methods for code smell detection. In this regard, we systematically evaluate the effectiveness of state-of-the-art PEFT methods on both small and large Language Models (LMs) for code smell detection. To begin, we constructed high-quality Java code smell datasets sourced from GitHub. We then fine-tuned four small LMs (SLMs) and five LLMs using various PEFT techniques, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, for code smell detection. Our comparison against full fine-tuning revealed that PEFT methods not only achieve comparable or better effectiveness but also consume less peak GPU memory. Our analysis further explored the performance of SLMs versus LLMs in the context of code smell detection. Generally, LLMs outperform SLMs on detecting Complex Conditional, Complex Method, and Feature Envy. However, SLMs demonstrate superior performance on detecting Data Class. We also investigated the impact of varying hyper-parameter settings and the availability of training data. Our results show that increasing the number of trainable parameters does not necessarily improve the effectiveness of PEFT methods for code smell detection. In contrast, we observed significant performance gains when the training dataset size was increased. Finally, our results show that LMs with PEFT methods outperform traditional heuristic-based detectors, advanced DL-based approaches, and general-purpose LLMs under In-Context Learning in the task of code smell detection. Specifically, we observe improvements in MCC ranging from 4.11% to 19.25% across the four types of code smells studied, demonstrating a more effective solution for this task. In conclusion, our work provides a comprehensive evaluation of PEFT methods on both small and large LMs for code smell detection. The results highlight the efficiency and effectiveness of PEFT methods in this domain, offering valuable insights for optimizing LMs for code smell detection. We suggest that the choice of PEFT method for code smell detection should be guided by the model used, data availability, and computational resources. Our replication package is available at (replicationPackage, ), enabling others to reproduce and extend our findings.
@@ 1. Introduction
Code smells are indicators of potential issues in code or design (fowler1999refactoring, ), which could adversely affect software quality, particularly in terms of maintainability (palomba2018maintainability, ), code readability (abbes2011empirical, ), and testability (chatzigeorgiou2010investigating, ). Detecting code smells can be used to identify software design issues that lead to potential problems in future development and maintenance (fowler1999refactoring, ; lanza2007object, ; moha2010decor, ). Manual identification of code smells can be time-consuming and costly (prokic2024prescriptive, ), prompting many researchers to explore automated methods for more efficient code smell detection. Existing automated code smell detection methods are typically heuristics-based (marinescu2004detection, ; morales2017on, ; moha2010decor, ). Traditional heuristics-based detectors for code smells rely on a set of manually curated features and thresholds to discriminate between smelly and non-smelly classes (azeem2019machine, ). While it has been shown that these static detectors provide reasonably accurate results, they are proven to have limitations that hinder their practical application (fernandes2016review, ; zhang2011code, ), particularly because their performance is highly dependent on the selection of thresholds. Besides, code smells identified by such heuristics-based methods can be subjectively interpreted by developers, leading to a significant gap between detection results and developers’ agreement (azeem2019machine, ; fontana2016antipattern, ; mantyla2006subjective, ; schumacher2010building, ).
To address these issues, many researchers have incorporated machine learning (ML) and deep learning (DL) algorithms to mitigate the performance instability in code smell detection that arises from the manual threshold-setting process in traditional detectors (fontana2013code, ; fontana2016comparing, ; sharma2021code, ; liu2021deep, ). By learning to discern between smelly and non-smelly code elements, ML and DL techniques offer a solution to the subjectivity problem inherent in heuristics-based methods (nucci2018detecting, ). However, despite these advancements, ML and DL methods still face significant challenges in achieving considerable performance for code smell detection (pereira2022code, ; alazba2023deep, ).
In recent years, in the field of Natural Language Processing (NLP), increased computational power, advanced machine learning techniques, and access to very large-scale data have led to the emergence of Large Language Models (LLMs). The exceptional performance of LLMs in NLP tasks (e.g., text classification, generation, and summarization) has garnered significant attention and interest from the Software Engineering (SE) domain. Researchers have begun employed LLMs, particularly code-specific LLMs like CodeBERT-Base (feng2020codebert, ) and CodeT5 (wang2021codet5, ), to tackle various SE tasks, such as code review (lu2023llamareviewer, ) and vulnerability repair (le2024study, ). However, directly using LLMs without any task-specific fine-tuning for code smell detection yields suboptimal results (mesbah2025leveraging, ; silva2024detecting, ; sadik2025benchmarking, ), as these models lack adaptation to the unique demands of the task.
Fine-tuning enables the adaptation of Language Models (LMs) to specific downstream tasks by updating model parameters, allowing them to better capture task-relevant features and improve performance. Among all the fine-tuning methods, Parameter-Efficient Fine-Tuning (PEFT) has emerged as a promising approach, offering an efficient alternative to full fine-tuning, particularly when computational resources are limited. PEFT methods have shown potential in adapting LLMs to various downstream tasks with fewer parameters and reduced resource requirements, thus expanding the practical applicability of LLMs in diverse SE contexts. Previous studies have explored the efficiency of PEFT methods for LMs on automated program repair (li2024comprehensive, ), code generation (liu2023empirical, ), code search (wang2023one, ), etc. However, to the best of our knowledge, no prior research has focused on improving the performance of LLMs for code smell detection, nor has any in-depth study been conducted analyzing the application of PEFT methods in this area. Additionally, identifying the code smells selected in this study requires both method- and class-level code understanding. While existing research on PEFT methods has predominantly focused on code-related tasks at the method level (li2024comprehensive, ; liu2023empirical, ; wang2023one, ; liu2024delving, ; weyssow2023exploring, ), a significant gap remains in assessing their effectiveness at the class level. By leveraging the task of code smell detection, our study aims to fill this gap by systematically exploring the capacities of LMs with PEFT methods in class-level code understanding. Furthermore, utilizing PEFT methods for code smell detection can facilitate just-in-time code smell detection. Traditional heuristic-based methods for code smell detection are often applied after the software systems have been completed (designitejava, ; pmd, ; marinescu2005iplasma, ). As a result, code smells are typically discovered at later stages of development, requiring refactoring that involves modifying all relevant parts of the system, thereby increasing the cost and complexity of addressing these code smells. In contrast, incorporating LLMs with PEFT techniques allows for real-time detection of code smells, enabling developers to identify and address potential code smells as they write code. For instance, upon completing a function, developers can immediately receive feedback on whether it contains a code smell. This proactive approach contributes to maintaining high code quality throughout the development life cycle, leading to more efficient and effective code maintenance and improvement (ardimento2021temporal, ; ardimento2021transfer, ).
To this end, we conduct a series of experiments on fine-tuning both small and large LMs with PEFT methods on the code smell detection task. In this paper, we define small LMs (SLMs) as those with fewer than 1B parameters, while LMs with 1B parameters or more are considered LLMs. Specifically, we construct high-quality Java code smell datasets comprising four types of widely studied and sophisticated code smells at both the method and class level, i.e., Complex Conditional (CC), Complex Method (CM), Feature Envy (FE), and Data Class (DC). These datasets are developed to address the lack of suitable existing resources, as most available datasets either do not conform to the required data formats or are based on outdated codebases. We then apply various PEFT methods, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, to fine-tune four popular SLMs and five cutting-edge LLMs with parameter sizes ranging from 125M to 7B. The effectiveness of these PEFT methods is assessed for code smell detection, and their performance is compared to that of full fine-tuning in terms of both effectiveness and peak GPU memory usage. Additionally, we investigate the performance of PEFT techniques under different hyper-parameter settings and evaluate their effectiveness in low-resource scenarios with varying numbers of training samples on the code smell detection task. Finally, we employ a heuristics-based static analysis tool alongside two state-of-the-art DL-based code smell detection tools to compare the performance of PEFT methods with existing approaches.
The main contributions of this work are as follows:
  • We mine Java repositories from GitHub and employ traditional heuristic-based detectors, along with a two-stage manual review, to create high-quality and balanced datasets for four types of code smells, i.e., Complex Conditional, Complex Method, Feature Envy, and Data Class.
  • We conduct a comprehensive evaluation of multiple PEFT methods on both small and large LMs for code smell detection. Our results indicate that LLMs, when fine-tuned, do not always outperform SLMs. Moreover, PEFT methods achieve better or comparable effectiveness compared to full fine-tuning. Overall, LMs with PEFT methods exhibit promising capabilities in method- and class-level code understanding.
  • We explore the performance of PEFT methods under different hyper-parameter settings and in low-resource scenarios. Our results reveal that adjusting hyper-parameters to increase trainable parameters does not necessarily lead to better effectiveness in detecting code smells. However, the performance of PEFT methods improves with more training samples.
  • We compare a heuristic-based detector, two best-in-class DL-based detection tools, and two general-purpose LLMs using In-Context Learning (ICL) against our PEFT-based method on code smell detection. Our results demonstrate that LMs with PEFT methods consistently outperform all baselines across multiple code smell types, highlighting the superiority of PEFT methods.
  • We provide actionable suggestions and insights for both developers and researchers for using PEFT methods on code smell detection.
@@
The remainder of this paper is organized as follows: Section 2 presents the background information. Section 3 outlines the research questions and the process of dataset construction, while the study results are provided in Section 4. Section 5 discusses the insights for practitioners. Potential threats to validity are addressed in Section 6, with related work reviewed in Section 7. Finally, Section 8 concludes the paper and suggests future directions for research.
2. Background
2.1. The Need of Applying Language Models on Code Smell Detection
@@
Existing heuristic-based code smell detectors rely on pre-defined rules and metrics, while ML- or DL-based approaches depend on algorithms to identify the optimal threshold for code smell classification. However, these tools often fail to consider the inherent context of the code, which is crucial for accurately detecting and understanding code smells (sharma2021code, ). For example, the Java method shown in Figure 1 contains a switch-case statement with seven cases. From the perspective of a Java developer, this method simply maps the numbers 1 through 7 to the days of the week, from Monday to Sunday. It is straight-forward and easy to understand, as all the case statements follow the same structure. However, when scanned by an existing heuristic-based code smell detector, such as DesigniteJava (designitejava, ), the tool flags the method with the Complex Method code smell. This decision is based solely on the method’s cyclomatic complexity, which exceeds a pre-defined threshold, without considering the method’s context.
In contrast, Language Models (LMs), such as ChatGPT (achiam2023gpt, ), have the advantage of considering the inherent context of the code (wang2021codet5, ; roziere2024codellama, ). The contextual understanding allows the LM to accurately determine that the method does not exhibit Complex Method, as it can perform a role similar to a skilled developer assessing the code. The LM can evaluate not only traditional metrics, such as cyclomatic complexity, but also analyze additional factors, including readability, modularity, and overall design. Therefore, compared to traditional heuristic-based static analysis tools and certain ML- or DL-based detectors, LMs have the potential to detect code smells more accurately by considering both structural and contextual factors, rather than relying solely on numerical thresholds or pre-defined rules.
Figure 1. An example of a Java method that is considered to contain Complex Method code smell by the heuristic-based tool
2.2. Full Fine-tuning

### v3

@@ Abstract.
Code smells are indicators of suboptimal design that negatively impact software quality. However, automated code smell detection remains a persistent challenge: heuristics-based tools suffer from high sensitivity to threshold selection and inherent subjectivity, while Machine Learning (ML) and Deep Learning (DL) models yield unsatisfactory performance. Although Large Language Models (LLMs) offer a promising solution to these limitations, their adaptation is impeded by the prohibitive cost of full fine-tuning and the lack of “LM-ready” benchmarks, as existing code smell datasets largely rely on software metrics or contain noisy, unverified labels. To address these gaps and enable a rigorous assessment of efficient LLM adaptation for code smell detection, we present a comprehensive study comprising two synergistic contributions. First, we constructed a high-quality, source-code-centric benchmark covering four non-trivial code smells—Complex Conditional, Complex Method, Feature Envy, and Data Class—and ensured its reliability through a meticulous two-stage manual review. Second, leveraging this benchmark, we systematically evaluated the effectiveness of four Parameter-Efficient Fine-Tuning (PEFT) methods (i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) across four Small LMs (SLMs) and five LLMs. We benchmarked these PEFT-tuned LMs against a broad suite of baselines, including traditional heuristics-based detectors (DesigniteJava and PMD), DL-based approaches (DeepSmells and AE-Dense), and cutting-edge general-purpose LLMs (GPT-4o-mini and DeepSeek-v3) with In-Context Learning. Our results demonstrate that PEFT methods match or exceed full fine-tuning performance while significantly reducing peak GPU memory usage, and outperform all baselines with MCC improvements ranging from 0.33% to 13.69%, particularly for some code smells. Furthermore, we provide actionable insights into the choice of PEFT method for code smell detection, which should be guided by the model used, data availability, and computational resources. Our replication package is available at (replicationPackage) to support reproducibility and future extensions.
@@ 1. Introduction
Code smells are indicators of underlying design or implementation issues that may degrade software quality (fowler1999refactoring), particularly in terms of maintainability (palomba2018maintainability), readability (abbes2011empirical), and testability (chatzigeorgiou2010investigating). Detecting code smells helps to uncover software design flaws that lead to potential problems in future development and maintenance (fowler1999refactoring; lanza2007object; moha2010decor). However, manual identification of code smells is often time-consuming and cost-prohibitive (prokic2024prescriptive), which has prompted extensive research into automated detection techniques.
Early work predominantly relied on heuristics-based detectors (marinescu2004detection; morales2017on; moha2010decor). These methods apply manually curated features and thresholds to discriminate between smelly and non-smelly instances (azeem2019machine). While effective in some settings, such static tools exhibit well-documented shortcomings (fernandes2016review; zhang2011code): their performance is highly sensitive to threshold selection, they offer limited semantic understanding, and their results often diverge from human judgment due to the inherent subjectivity of code smells (azeem2019machine; fontana2016antipattern; mantyla2006subjective; schumacher2010building).
To overcome these limitations, researchers have increasingly incorporated Machine Learning (ML) and Deep Learning (DL) algorithms for code smell detection. These approaches aim to mitigate the performance instability inherent in the manual threshold-setting of heuristics-based detectors (fontana2013code; fontana2016comparing; sharma2021code; liu2021deep). By reducing reliance on rigid thresholds, ML and DL models can partially learn semantic cues (nucci2018detecting). Nonetheless, recent studies indicate that these models still struggle to achieve considerable performance (pereira2022code; alazba2023deep), as they frequently fail to capture the deeper semantic and structural relationships within source code.
Recently, Large Language Models (LLMs) have emerged as a transformative force in Software Engineering (SE), achieving state-of-the-art performance across a wide range of tasks (lu2023llamareviewer; le2024study). Despite their impressive general capabilities, directly employing off-the-shelf LLMs for code smell detection without any task-specific fine-tuning yields suboptimal results (mesbah2025leveraging; silva2024detecting; sadik2025benchmarking). While effective adaptation is essential, two primary challenges currently impede progress:
Challenge 1: Lack of High-Quality, LM-Ready Datasets. Most existing code smell datasets comprise calculated software metrics rather than the raw source code required by LMs. Moreover, they typically focus on relatively simple smells that do not demand complex semantic reasoning. Furthermore, datasets that do include code often rely on noisy labels generated directly by heuristics-based detectors without human verification, leading to the “Oracle Problem”. This incompatibility renders existing benchmarks unsuitable for the effective training or evaluation of LMs in code smell detection. Consequently, without high-quality, source-code-centric benchmarks covering non-trivial code smells, a fair investigation into the effectiveness of LMs for code smell detection is infeasible.
Challenge 2: Unclear Effectiveness of LM Adaptation Methods. Although Parameter-Efficient Fine-Tuning (PEFT) methods have shown promise across various SE tasks by enabling task-specific adaptation with only a small number of trainable parameters (wang2023one; liu2023empirical; he2022towards; liu2024delving; weyssow2023exploring; li2024comprehensive), their effectiveness for code smell detection remains largely unexplored. It is still unclear whether PEFT approaches can match the performance of full fine-tuning or outperform existing baselines in code smell detection.
Investigating PEFT methods in the context of code smell detection also fills a broader research gap. Prior studies on PEFT have primarily concentrated on method-level tasks (li2024comprehensive; liu2023empirical; wang2023one; liu2024delving; weyssow2023exploring), leaving their effectiveness in class-level contexts largely unexplored. Our study addresses this gap by exploring the capabilities of PEFT-adapted LMs in capturing class-level semantics through the detection of code smells, such as Feature Envy and Data Class, which inherently necessitate class-level code understanding. Furthermore, utilizing PEFT techniques facilitates just-in-time code smell detection. Traditional heuristics-based detectors are often applied post-development (designitejava; pmd; marinescu2005iplasma), discovering issues only after software systems are fully implemented and thereby increasing refactoring overhead. In contrast, PEFT-tuned LMs enable real-time identification of code smells, allowing developers to receive immediate feedback upon completing a code block. This proactive approach supports high code quality throughout the development lifecycle, leading to more efficient software maintenance (ardimento2021temporal; ardimento2021transfer).
To this end, we systematically evaluate the effectiveness of state-of-the-art PEFT methods on both Small LMs (SLMs, <<1B parameters) and Large LMs (LLMs, $\geq$1B parameters) for code smell detection. To facilitate this evaluation and resolve the aforementioned data bottleneck, we first curated a high-quality Java code smell dataset covering four types of widely studied and sophisticated code smells at both method and class levels: Complex Conditional (CC), Complex Method (CM), Feature Envy (FE), and Data Class (DC). Crucially, we subjected this dataset to a rigorous two-stage manual review to guarantee its reliability. With a robust benchmark established, we then addressed the adaptation challenge by fine-tuning four SLMs and five LLMs using various PEFT techniques (i.e., prompt tuning, prefix tuning, LoRA, and (IA)3). We compared these PEFT-tuned LMs against a comprehensive suite of baselines, encompassing heuristics-based detectors, DL-based approaches, and general purpose LLMs with In-Context Learning (ICL).
The main contributions of this work are summarized as follows:
  • LM-Ready Benchmark for Code Smell Detection: We construct a high-quality, source-code-centric dataset covering four non-trivial code smells—Complex Conditional, Complex Method, Feature Envy, and Data Class. By mining recent Java repositories, applying heuristics-based detectors, and performing a stringent two-stage manual review, we address the limitations of existing code smell datasets (e.g, reliance on code metrics, focusing on simple smells, and unverified labels), ensuring high data quality and a balanced distribution. This establishes a reliable ground truth necessary for training and evaluating LMs for code smell detection, serving as the foundation for our subsequent experiments.
  • Systematic Evaluation of PEFT Methods: Leveraging our curated dataset, we conduct the first extensive evaluation of PEFT methods for code smell detection on both SLMs and LLMs. We compare multiple PEFT techniques, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, against full fine-tuning. Our analysis comprehensively assesses their effectiveness, peak GPU memory usage, and performance under different hyper-parameter configurations and low-resource scenarios.
  • Comprehensive Comparison with SOTA Baselines: To validate the utility of PEFT-tuned LMs for code smell detection, we benchmark these techniques against a wide range of baselines. These baselines span traditional heuristics-based tools, advanced DL-based detectors, and general-purpose LLMs with In-Context Learning under zero-shot, few-shot, and criteria-informed settings. Our results demonstrate that PEFT methods consistently achieve comparable or superior effectiveness while reducing computational cost, highlighting their practical value for the code smell detection.
  • Practical Insights for Developers and Researchers: Based on our extensive experiments, we distill actionable recommendations for applying PEFT methods to code smell detection. We offer specific guidance on model selection, hyper-parameter tuning strategies, and the trade-offs between efficiency and performance, providing a useful reference for future research and practice in this domain.
@@
The remainder of this paper is organized as follows: Section 2 introduces the background and motivation. Section 3 details the dataset construction process, while Section 4 outlines the research questions and experimental setup. The study results are presented in Section 5, followed by a discussion in Section 6. Threats to validity are addressed in Section 7, with the related work reviewed in Section 8. Finally, Section 9 concludes the paper.
2. Background and Motivation
2.1. The Case for Applying Language Models on Code Smell Detection
@@
Existing heuristics-based code smell detectors rely on predefined rules and metrics, while ML- or DL-based approaches depend on algorithms to identify the optimal threshold for code smell classification. However, these tools often overlook the semantic and contextual information of code, which is crucial for accurately detecting and understanding code smells in real-world software development (sharma2021code). For instance, the Java method shown in Figure 1 contains a switch-case statement with seven cases. From the perspective of a Java developer, this method simply maps the numbers 1 through 7 to the days of the week. It is straightforward and easy to understand, as all the case statements follow the same structure. While the method could be refactored to reduce repetition, it has a single responsibility, and its logic is immediately clear to a reader. Therefore, the method does not exhibit a Complex Method (CM) code smell in practice. However, when scanned by the heuristics-based code smell detector, DesigniteJava (designitejava), the method is flagged as CM solely because its cyclomatic complexity exceeds a predefined threshold, while its semantic clarity is ignored.
Figure 1. An example of Complex Method (CM) code smell detection by DesigniteJava, ChatGPT, and human developers
public static String format2Duration(long ms) {
long days = MILLISECONDS.toDays(ms);
long hours = MILLISECONDS.toDurationHours(ms);
long minutes = MILLISECONDS.toDurationMinutes(ms);
long seconds = MILLISECONDS.toDurationSeconds(ms);
if (days == 0 && hours == 0 && minutes == 0 && seconds == 0) {
seconds = 1;
StringBuilder strBuilder = new StringBuilder();
if (days > 0) strBuilder.append(days).append("d ");
if (hours > 0) strBuilder.append(hours).append("h ");
if (minutes > 0) strBuilder.append(minutes).append("m ");
if (seconds > 0) strBuilder.append(seconds).append("s");
return strBuilder.toString();
(a) False positive: identified as CC-positive by DesigniteJava but clean labeled by developers
public boolean onActionItemClicked(ActionMode mode, MenuItem menuItem) {
int id = menuItem.getItemId();
if (id == R.id.delete_entry) {
confirmDelete(mode);
return true;
} else if (id == R.id.select_all) {
if (mSelected.size() >= mAdapter.getCount())
mode.finish();
else {
for (int i = 0; i < mAdapter.getCount(); i++) {
if (!mListView.isItemChecked(i))
mListView.setItemChecked(i, true);
return true;
} else {
return false;
(b) False negative: identified as clean by DesigniteJava but contains developer-labeled CC code smell
Figure 2. Examples of Complex Conditional (CC) code smell detection by DesigniteJava, highlighting a false positive (a) and a false negative (b)
Figure 2 illustrates additional limitations in rule-based detection for Complex Conditional (CC) code smells. Example (a) is identified as CC-positive by DesigniteJava but is considered clean by human developers: the conditional if (days == 0 && hours == 0 && minutes == 0 && seconds == 0) exceeds the operand threshold predefined by DesigniteJava, yet the logic is simple and readable, resulting in a false positive. Example (b) is labeled as clean by DesigniteJava but contains a developer-identified CC: nested loops and interacting conditionals make the overall logic difficult to follow, even though no single conditional exceeds structural thresholds, resulting in a false negative.
In contrast, Language Models (LMs) such as ChatGPT (achiam2023gpt) can leverage semantic and contextual understanding (wang2021codet5; roziere2024codellama), enabling detection of both CM and CC code smells consistent with human judgment. For CM, LMs can recognize that a method with high cyclomatic complexity or many lines of code may still be clear and maintainable if its logic is transparent and responsibilities are well-defined. For CC, LMs can evaluate not only structural metrics, such as nesting depth or the number of boolean operands, but also semantic aspects such as readability, logical flow, and how the conditions interact, allowing them to correctly identify both false positives and false negatives that rule-based tools might misclassify. By integrating structural and semantic reasoning, LMs offer a more accurate and nuanced approach than traditional heuristics or purely threshold-based ML or DL tools. However, applying LMs on code smell detection effectively requires high-quality data and efficient training strategies, which motivates our work.
2.2. The Gap in Code Smell Detection Datasets
Despite the growing use of Large Language Models (LLMs) in software engineering research, existing code smell datasets remain a major bottleneck for effective model training. We identify three critical limitations that make current public datasets inadequate for fine-tuning LMs for code smell detection.
1. Predominance of Metric-Based Data: Traditional code smell detection treated the task as tabular classification, leading to widely used datasets containing only software metrics rather than actual source code (fontana2016comparing; binh2022mlcodesmell; madeyski2023detecting). While suitable for training classical classifiers like Random Forests or Logistic Regression, these metric-based datasets are incompatible with LMs, which require tokenized source code to capture semantic and syntactic patterns.
2. Limited Scope of Code Smell Types: Existing datasets often lack diversity in the types of smells covered (fontana2016comparing; madeyski2020mlcq; fontana2017code; liu2021deep; moha2009decor). There is a disproportionate focus on easily detectable, metric-based smells (e.g., Long Method). Conversely, there is a scarcity of high-quality samples for structurally intricate smells, such as Complex Conditional and Complex Method, which necessitate semantic reasoning beyond simple metric thresholds.
3. The “Oracle Problem” and Label Reliability: A significant number of large-scale datasets rely on heuristics-based detectors (e.g., PMD, DECOR, and iPlasma) as the “oracle” to automatically generate labels, typically bypassing manual verification (moha2009decor; lenarduzzi2019technical; sae2016context; yang2015classification; khomh2012exploratory). However, the inherent inaccuracies of these tools introduce noise into the ground truth (palomba2018diffuseness; de2017towards). Training an LM on such datasets risks merely teaching the model to emulate the flaws and biases of heuristics-based detectors rather than identifying genuine design issues.
In summary, current code smell datasets are not “LM-ready”. To effectively fine-tune an LM for code smell detection, there is an urgent need for a dataset that provides raw, compilable source code, covers complex semantic smells, and relies on high-quality, human-validated labels. This motivates the construction of our proposed dataset.
2.3. Full Fine-Tuning and Parameter-Efficient Fine-Tuning (PEFT)
2.3.1. Full Fine-Tuning